Predict which bond has the most covalent character.

Group of answer choices

C—S

Si—N

O — F

Na—Br

B—Cl
Answer #1

Solution: Covalent character bond is c-s Si-n o- Na Br B-ce F 2.5-2.5 =o) 13-18 = 12) (4-35 = 0.5 12.8 - 0.g= 1.9) 13.0-2.0=